The table below shows the supply and demand conditions for a band that will play music on the streets when requested. Qs1 is the quantity supplied without social costs. Qs2 is the quantity supplied with social costs.Price Qs1 Qs2 $20 o 10 8 1 9 7 $18 $15 2.5 7.5 5.5 4 6 4 $12 $10 5 5 3 $5 7.5 2.5 0.5

Answer the following:

a) What is the negative externality in this situation?

b) Identify the equilibrium price and quantity when we account only for private costs.

c) Identify the equilibrium price and quantity when we account for social costs.

d) How does accounting for the externality affect the equilibrium price and quantity
Answer #1

(a)

Negative externality = Vertical distance between QS1 and Qs2 = 10 - 8 = 9 - 7 = ..... = 2.5 - 0.5 = $2

(b)

With only private costs, QD = Qs1 = 5 units with Price = $10

(c)

Considering externality, QD = Qs2 = 4 units with Price = $12

(c)

When externality is considered, Price is higher and quantity is lower.